Two cards are drawn, with replacement, from a standard 52-card deck. Find the expected number of diamonds
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

Two cards are drawn, with replacement, from a standard 52-card deck. Find the expected number of diamonds
----
On each draw the expected # of diamonds is (1/4)(13) = 3.25
---
Cheers,
Stan H.
------------------------------------

Answer by Edwin McCravy(20054)



The other tutor's answer is incorrect.

The possibilities are 

1. 0 diamonds, that is, a non-diamond the first time and 
   a non-diamond the second time.
  
2. 1 diamond, that is, a diamond the first time and a non-diamond 
   the second time, or a non-diamond the first time and a diamond 
   the second time.
 
3. 2 diamonds, that is, a diamond the first time and a diamond
   the second time.


Remember that "AND" usually indicates that you are to multiply 
probabilities, while "OR" indicates that you are to add them.

The probability of drawing a diamond is 1/4 and the probability of 
drawing a non-diamond is 3/4

1. Find the probability of 0 diamonds, (two non-diamonds)
   P(non-diamond first AND non-diamond second) =
   P(non-diamond first)*P(non-diamond second)
   (3/4)(3/4) = 9/16 
   
2. Find the probability of exactly 1 diamond (and 1 non-diamond)
   P[(diamond first AND non-diamond second) OR (non-diamond first
   AND diamond second) =
   P(diamond first)*P(non-diamond second) + 
                P(non-diamond first)*P(diamond second) =
   (1/4)(3/4) + (3/4)(1/4) = 3/16 + 3/16 = 6/16 = 3/8 

3. Find the probability of 2 diamonds:
   P(diamond first AND diamond second) =
   P(diamond first)*P(diamond second) =
   (1/4)(1/4) = 1/16
   
Next we make a discrete probability distribution function:

Number of   
diamonds    Probability
   x           P(x)
------------------------------------
   0           9/16
   1            3/8 
   2           1/16
------------------------------------
                 1   


Notice that the sum of the probabilities of all possible events is
9/16 + 3/8 + 1/16 = 9/16 + 6/16 + 1/16 = 16/16 = 1,
which you should always check. 

To find the expectation of the variable x:

E(x) = ∑[x*P(x)] = 0*(9/16) + 1*(3/8) + 2*(1/16) = 

0 + 3/8 + 2/16 = 3/8 + 1/8 = 4/8 = 1/2

So the expectation is 1/2.

[Note: "Expectation" means what you would expect to average if
you repeated the experiment every day for a long period of time
and took the average of the number of diamonds you drew. It does
NOT mean that you would ever "expect" to get half a diamond. 
However if you drew two cards without replacement every day for
365 days, you might get say 181 diamonds, and you would have
averaged 181/365 or .49589 diamonds per day, which is very near .5
or 1/2.]

Edwin